OpenStudy (anonymous):

Create your own unique literal equation and solve for one of the variables. Show your work. Then, using complete sentences, explain how you solved for the variable you chose.. Need some help again /:

OpenStudy (anonymous):

a literal equation, i think... has more than one variable

OpenStudy (anonymous):

so an equation like..... 2a(3b)=(1/c)

OpenStudy (anonymous):

is literal

OpenStudy (anonymous):

But, is that how you do it? I'm seriously terrible in math /: Sorry if i sound stupid.

OpenStudy (anonymous):

if you want to solve for (and in this question, i dont think it matter which variable you solve for) b.... you could multiply the coeeficients to get 6ab=1/c

OpenStudy (anonymous):

then, divide both sides by 61... you get b=1/6ac

OpenStudy (anonymous):

im sorry, divide both sides by 6a

OpenStudy (anonymous):

does this make sense? are you caught up on something?

OpenStudy (anonymous):

I'm doing flvs for algebra 2, cause I failed it so I have to make up the credit /: But, it kinda makes sense I'm just really bad at math..

OpenStudy (anonymous):

a little practice and you can be great at math... just remember a literal equation has more than 1 variable... take make a unique one you can literally just throw together variables and coefficients together, use your normal algebra skills to get a variable by itself on one side of the equation

OpenStudy (anonymous):

good luck :)

OpenStudy (anonymous):

Thanks (: